The cost of a class picnic was $42. If 14 more students had come, the cost per student would have been 20 cents less. How much is the cost per student?
Answer by stanbon(48558)

The cost of a class picnic was $42. If 14 more students had come, the cost per student would have been 20 cents less. How much is the cost per student?
-------------------------------------------------
Let original number of students is "x".
---------
Original cost per student = 42/x dollars
---------------------------------------------
When 14 students are added the cost per student = 42/(x+14)
----------------------------------------------
Equation:
original cost per student - new cost per student = 1/5 dollar
42/x - 42/(x+14) = 1/5
----
Multiply thru by 5x(x+14) to get:
42*5(x+14) - 42*5x = x(x+14)
210x + 210*14 - 210x = x^2 + 14x
x^2 + 14x - 2940 = 0
-------
x = [-14 +- sqrt(14^2 - 4*1*-2940)]/2
x = 47.67
---------
Note: The number should be a positive whole number.
Check you data. Are you sure is was 14 students
and not 4 that were added?
Cheers,
Stan H.